How many grams of O2 will completely react with 49.0 g PH3?



1
Expert's answer
2021-10-12T02:04:44-0400

2PH3 + 4O2 P2O5 + 3H2O

M(PH3) = 40 g/mol

M(O2) = 32 g/mol

n = m/M

n(PH3) = 49/40 = 1.225 mole

n(O2) = 2n(PH3) "= 2\times 1.225 = 2.45 \; mol"

"m = n\times M"

m(O2) "= 2.45 \times 32 = 78.4 \;g"